Apply the EOQ model to the following quantity discount situation in which D = 400 units per year, Co = $44, and the annual holding cost rate is 20%.

Discount
Category

Order Size
Discount
(%)

Unit Cost
1   0 to 99 0 $10.00  
2   100 or more 3 $9.70  

What order quantity do you recommend? If required, round your answer to the nearest whole number.
Answer #1

EOQ Quantity = (2*Annual Demand*Ordering Cost per order/Carrying cost per Unit)^1/2

= (2*400*44/2)

= 132.66 units

Number of orders = 400/133 = 3

Since the discount is also available at EOQ, 133 units should be purchased at a time

3 orders of 133, 133 and 134 units